# La RoSA: Enhancing LLM Efficiency via Layerwise Rotated Sparse Activation

### v1

@@ Appendix A Theoretical Analysis
This section provides a detailed explanation of how LaRoSA utilizes orthogonal transformations and top-k sparsification, and why it achieves consistently lower empirical errors compared to TEAL (Liu et al., 2024a).
TEAL derives the theoretical errors from magnitude-based sparsification under a restrictive assumption that weights and activations are independent and identically Gaussian distributed. However, we found that empirical errors are generally larger than theoretical errors, as shown in  Figure 5, which may be due to the incorrect assumption of a Gaussian distribution.
@@
$\frac{\mathbb{E}\|{\bm{y}}-{\bm{y}}^{S}\|_{2}}{\mathbb{E}\|{\bm{y}}\|_{2}}=\sqrt{1-\frac{k}{D_{\text{in}}}-2\Phi^{-1}\left(1-\frac{k}{2D_{\text{in}}}\right)\varphi\left(\Phi^{-1}\left(1-\frac{k}{2D_{\text{in}}}\right)\right)}$ (11)
@@
$\mathbb{E}[\hat{{\bm{y}}}_{j}\hat{{\bm{y}}}_{k}]$ $=\mathbb{E}\left[\sum_{i_{1}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{1}}-S_{k}(\tilde{{\bm{x}}}_{i_{1}}))\tilde{{\mathbf{W}}}_{ji_{1}}\sum_{i_{2}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{2}}-S_{k}(\tilde{{\bm{x}}}_{i_{2}}))\tilde{{\mathbf{W}}}_{ki_{2}}\right]$
$=\mathbb{E}\left[\sum_{i_{1}=1}^{D_{\text{in}}}\sum_{i_{2}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{1}}-S_{k}(\tilde{{\bm{x}}}_{i_{1}}))(\tilde{{\bm{x}}}_{i_{2}}-S_{k}(\tilde{{\bm{x}}}_{i_{2}}))\tilde{{\mathbf{W}}}_{ji_{1}}\tilde{{\mathbf{W}}}_{ki_{2}}\right]$
@@
Also for any $i$, according to Lemma A.1 in  (Liu et al., 2024a), we have :
@@
$=\sigma_{w}^{2}\sum_{j=1}^{D_{\text{in}}}\mathbb{E}[\tilde{{\bm{x}}}_{j}^{2}\big{|}\tilde{{\bm{x}}}_{j}\notin Top_{k}(\tilde{{\bm{x}}})]$
@@
On the other hand, since $\tilde{{\bm{x}}}_{i}$ are Gaussian distributed, $11-\frac{k}{D_{\text{in}}}=\mathbb{P}(\lvert\tilde{{\bm{x}}}_{i}\rvert\leq t_{k})=\mathbb{P}(\lvert z\rvert\leq\frac{t_{k}}{\sigma_{x}})=2\Phi(\frac{t_{k}}{\sigma_{x}})-$ 1, where $z$ is a standard Gaussian random variable. It follows that:
@@
as the unsparsified case. Equation 15 and Equation 16 lead to the final result. ∎
@@
As illustrated in  Figure 5, the improved Gaussian distribution assumption results in LaRoSA having consistent lower empirical errors compared to TEAL, bringing our proposed method closer to the theoretical errors. We employ magnitude-based top-K method instead of pruning the last few elements after rotation like SliceGPT (Ashkboos et al., 2024a), which ensures the validity of the aforementioned theoretical error analysis. Our experiments across multiple models demonstrate that, after rotating into the principal component space, most small-magnitude elements of activation are indeed concentrated in the last few dimensions.
@@ Appendix B Grid Search Optimized Top-K
After obtaining the optimal coefficients $\alpha$ for ${\bm{h}}_{1},{\bm{h}}_{2},{\bm{h}}_{3},$ and ${\bm{h}}_{4}$, we apply the same allocation weights across all layers. The experimental results in Table 10 demonstrated that our sparsity coefficients search strategy effectively improves the performance of the sparsified model. In Table 11, we provide a detailed presentation of the sparsity coefficients used across various models.
@@
[…]
@@
10:        if current_PPL $<$ best_PPL then
[…]
@@
19:  $\alpha_{2}^{*}\leftarrow 4-3*\alpha_{1}^{*}$
20:  $\alpha_{4}^{*}\leftarrow(2+M-2*\alpha_{3}^{*})/M$
@@ Appendix C Full Experimental Results
LLaMA2-7B 29.26(1.00$\times$) 31.46(1.07$\times$) 29.26(1.00$\times$) 31.46(1.07$\times$)
LLaMA2-70B 292.44(1.00$\times$) 314.43(1.07$\times$) 292.45(1.00$\times$) 314.44(1.07$\times$)
LLaMA3-8B 32.94(1.00$\times$) 35.14(1.06$\times$) 32.94(1.00$\times$) 35.14(1.06$\times$)
LLaMA3-70B 295.67(1.00$\times$) 317.66(1.07$\times$) 295.68(1.00$\times$) 317.67(1.07$\times$)
Mistral-7B 31.34(1.00$\times$) 33.53(1.07$\times$) 31.34(1.00$\times$) 33.54(1.07$\times$)
Qwen2.5-7B 30.64(1.00$\times$) 32.11(1.05$\times$) 30.64(1.00$\times$) 32.12(1.05$\times$)
Qwen2.5-72B 303.69(1.00$\times$) 32.57(1.07$\times$) 303.69(1.00$\times$) 325.68(1.07$\times$)

### v2

@@ Appendix A Theoretical Analysis
This section provides a detailed explanation of how LaRoSA utilizes orthogonal transformations and top-k sparsification, and why it achieves consistently lower empirical errors compared to TEAL (liu2024teal).
TEAL derives the theoretical errors from magnitude-based sparsification under a restrictive assumption that weights and activations are independent and identically Gaussian distributed. However, we found that empirical errors are generally larger than theoretical errors, as shown in  Figure˜5, which may be due to the incorrect assumption of a Gaussian distribution.
@@
$\frac{\mathbb{E}\|{\bm{y}}-{\bm{y}}^{S}\|_{2}}{\mathbb{E}\|{\bm{y}}\|_{2}}=\sqrt{1\hskip-3.0pt-\hskip-3.0pt\frac{k}{D_{\text{in}}}\hskip-3.0pt-\hskip-3.0pt2\Phi^{-1}\left(1\hskip-3.0pt-\hskip-3.0pt\frac{k}{2D_{\text{in}}}\right)\varphi\left(\Phi^{-1}\left(1\hskip-3.0pt-\hskip-3.0pt\frac{k}{2D_{\text{in}}}\right)\right)}$ (11)
@@
$\mathbb{E}[\hat{{\bm{y}}}_{j}\hat{{\bm{y}}}_{k}]$ $=\mathbb{E}\left[\sum_{i_{1}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{1}}\hskip-3.0pt-\hskip-3.0ptS_{k}(\tilde{{\bm{x}}}_{i_{1}}))\tilde{{\mathbf{W}}}_{ji_{1}}\sum_{i_{2}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{2}}\hskip-3.0pt-\hskip-3.0ptS_{k}(\tilde{{\bm{x}}}_{i_{2}}))\tilde{{\mathbf{W}}}_{ki_{2}}\right]$
$=\mathbb{E}\left[\sum_{i_{1}=1}^{D_{\text{in}}}\sum_{i_{2}=1}^{D_{\text{in}}}(\tilde{{\bm{x}}}_{i_{1}}\hskip-3.0pt-\hskip-3.0ptS_{k}(\tilde{{\bm{x}}}_{i_{1}}))(\tilde{{\bm{x}}}_{i_{2}}\hskip-3.0pt-\hskip-3.0ptS_{k}(\tilde{{\bm{x}}}_{i_{2}}))\tilde{{\mathbf{W}}}_{ji_{1}}\tilde{{\mathbf{W}}}_{ki_{2}}\right]$
@@
Also for any $i$, according to Lemma A.1 in  (liu2024teal), we have :
@@
[…]
@@
On the other hand, since $\tilde{{\bm{x}}}_{i}$ are Gaussian distributed, $1-\frac{k}{D_{\text{in}}}=\mathbb{P}(\lvert\tilde{{\bm{x}}}_{i}\rvert\leq t_{k})=\mathbb{P}(\lvert z\rvert\leq\frac{t_{k}}{\sigma_{x}})=2\Phi(\frac{t_{k}}{\sigma_{x}})-1$, where $z$ is a standard Gaussian random variable. It follows that:
@@
as the unsparsified case. Equation˜15 and Equation˜16 lead to the final result. ∎
@@
As illustrated in  Figure˜5, the improved Gaussian distribution assumption results in LaRoSA having consistent lower empirical errors compared to TEAL, bringing our proposed method closer to the theoretical errors. We employ magnitude-based top-K method instead of pruning the last few elements after rotation like SliceGPT (ashkboos2024slicegpt), which ensures the validity of the aforementioned theoretical error analysis. Our experiments across multiple models demonstrate that, after rotating into the principal component space, most small-magnitude elements of activation are indeed concentrated in the last few dimensions.
@@ Appendix B Grid Search Optimized Top-K
After obtaining the optimal coefficients $\alpha$ for ${\bm{h}}_{1},{\bm{h}}_{2},{\bm{h}}_{3},$ and ${\bm{h}}_{4}$, we apply the same allocation weights across all layers. The experimental results in Table˜10 demonstrated that our sparsity coefficients search strategy effectively improves the performance of the sparsified model. In Table˜11, we provide a detailed presentation of the sparsity coefficients used across various models.
@@
$\text{TopK+GS}_{60\%}$ 0.90 1.30 0.90 1.06 8.98
$\text{TopK+GS}_{60\%}$ 0.85 1.45 0.70 1.17 8.94
$\text{TopK+GS}_{60\%}$ 0.80 1.60 0.80 1.12 8.86
$\text{TopK+GS}_{60\%}$ 0.75 1.75 0.75 1.14 8.93
@@
10:   if current_PPL << best_PPL then
11:    $\text{best\_PPL}\leftarrow\text{current\_PPL}$
@@
[…]
@@ Appendix C Full Experimental Results
[…]